# Tony Lindeberg

### v7

@@ Abstract
To firmly determine if the underlying working hypothesis, regarding the receptive fields spanning a variability in the degree of elongation, would truly hold for the receptive fields in the primary visual cortex of higher mammals, we formulate a set of testable predictions, that can be used for investigate this property experimentally, and, if applicable, then also characterize if such a variability would, in a structured way, be related to the pinwheel structure in the visual cortex.
@@
Receptive field Elongation Affine covariance Orientation selectivity Gaussian derivative Quasi quadrature Simple cell Complex cell Pinwheel Vision Theoretical neuroscience
@@
Figure 2: Variability in the elongation of affine Gaussian derivative receptive fields (for image orientation $\varphi=0$), with the scale parameter ratio $\kappa=\sigma_{2}/\sigma_{1}$ increasing from $1$ to $4\sqrt{2}$ according to a logarithmic distribution, from left to right, with the vertical scale parameter kept constant $\sigma_{2}=4$ and with the horizontal scale parameter being the smaller $\sigma_{1}\leq\sigma_{2}$. (first row) First-order directional derivatives of affine Gaussian kernels according to (5) for $m=1$. (second row) Second-order directional derivatives of affine Gaussian kernels according to (5) for $m=2$. (third row) Second-order directional derivatives of affine Gaussian kernels according to (5) for $m=3$. (fourth row) Second-order directional derivatives of affine Gaussian kernels according to (5) for $m=4$. When the image structures are subject to foreshortening transformations, because of varying slant angles between the local surface normals and the viewing direction, the shapes of the spatial components of the receptive fields need to be adapted accordingly, to achieve affine covariance in the sense that the receptive field responses should be possibly to be appropriately matched under different viewing conditions. (Horizontal axes: image coordinate $x_{1}\in[-16,16]$. Vertical axes: image coordinate $x_{2}\in[-16,16]$.)
@@ 1 Introduction
When observing objects and events in our natural environment, the image structures in the visual stimuli will be subject to substantial variabilities caused by the natural image transformations. Specifically, if observing a smooth local surface patch from different viewing directions and viewing distances, this variability can, to first order of approximation, be approximated by local affine transformations (the derivative of the projective mappings between the different views). Within the 4-D variability of general centered 2-D affine transformations, a 1-D variability in the slant angle of the surface normal relative to the viewing direction does, in terms of covariance properties, correspond to a variability in the elongation of the receptive fields, if we would like the responses to be possible to perfectly match under such variabilities (see Figures 1 and 2 for illustrations and Section 2.1 for a complementary theoretical background).
In Lindeberg (2021; 2023; 2024b), we outlined a framework for how covariance properties with respect to geometric image transformations may constitute a fundamental constraint for the receptive fields in the primary visual cortex of higher mammals, to enable the visual computations to be robust under the variabilities in the image structures generated by the natural image transformations. According to the presented theory, based on axiomatically determined receptive field shapes derived from symmetry properties that reflect structural properties of the environment, in combination with additional constraints to guarantee consistency between image representations over multiple spatial and temporal scales, the population of receptive fields in the primary visual cortex ought to, according to this theory, obey covariance properties with respect to spatial affine transformations and Galilean transformations.
@@
With the lack of such neurophysiological data regarding full receptive field recordings, one could, however, aim to instead obtain indirect cues regarding a possible variability in the degree of elongation of the receptive fields, by making use of the recordings of the orientation selectivity of visual neurons by Nauhaus et al. (2008), which show a substantial variability regarding broad vs. sharp tuning of the receptive fields in the primary visual cortex, as well as by Goris et al. (2015), who report comparably uniform distributions of the degree of orientation selectivity of simple and complex cells, in terms of histograms of the resultant of the orientation selectivity curves.
@@
In a companion paper (Lindeberg 2024a), we established a strong direct link between the orientation selectivity and the elongation of the receptive fields according to the idealized generalized Gaussian derivative model for visual receptive fields (as will be summarized in Section 2.2). If we would assume that the generalized Gaussian derivative model would constitute a sufficiently valid model for the population of simple and complex cells in the primary visual cortex, then we could logically infer possible indirect support for the working hypothesis, in that the observed variability in the orientation selectivity of the receptive fields would correspond to a variability in the degree of elongation of the receptive fields.
1.1 The hypothesis about affine covariant receptive fields
A general motivation for the wider underlying working hypothesis about affine covariance is that, if the population of receptive fields would support affine covariance in the primary visual cortex, or sufficiently good approximations thereof, then such an ability would support the possibility of computing affine invariant image representations at higher levels in the visual hierarchy (Lindeberg 2013b), or more realistically sufficiently good approximations thereof, over restricted subspaces or subdomains of the most general forms of full variability under spatial affine transformations of the visual stimuli.
When we look at a slanted surface in the world, we can get a robust and stable perception of its surface texture under substantial variations of the slant angle. This robustness of the visual perception system under stretchings of image patterns that correspond to non-uniform scaling transformations (the perspective effects on a slanted surface patch can, to first-order of approximation, be modelled as a stretching of the image pattern along the tilt direction in image space, complemented with a uniform scaling transformation). If the visual receptive fields would span a variability under such spatial stretching transformations, then such a variability would precisely correspond to a variability in the anisotropy, or the degree of elongation, of the receptive fields.
1.2 Variability over the degree of elongation of receptive fields
The overall theme of this study is to, based on a theoretical analysis of a relationship between the orientation selectivity and the degree of anisotropy or degree of elongation of the receptive fields in the primary visual cortex, in combination with existing neurophysiological results concerning variabilities in the orientation selectivity of the receptive fields in relation to the pinwheel structure of higher mammals, address the question of whether the receptive fields in the primary visual cortex could be regarded as spanning a variability over the elongation of the receptive fields, to support covariant image measurements under such variabilities, or, at least, sufficiently good approximations thereof.
In particular, if we could assume that the spatial components of the receptive fields could be well modelled by affine Gaussian derivatives, then the property of affine covariance implies that there should be visual receptive fields for different degrees of anisotropy present in the visual system. If we further combine the theoretical results used in this paper, which state that the degree of orientation selectivity is strongly dependent on the anisotropy or elongation of the receptive fields, with the biological results established by Nauhaus et al. (2008), which show that the degree of orientation selectivity for neurons in the primary visual cortex varies both strongly and in relationship with the position on the cortical surface in relation to the pinwheel structure. Then, these results together are fully consistent with the hypothesis that the visual receptive fields in the primary visual cortex should span a variability in their anisotropy, thus consistent with the hypothesis that the receptive fields in the pinwheel structure should span at least one more degree of freedom in the affine group, beyond mere rotations (as already established in previous neurophysiological measurements regarding the pinwheel structure of the oriented receptive fields in the primary visual cortex of higher mammals).
@@ 1.2 Variability over the degree of elongation of receptive fields
We will also, more generally, use predictions from the presented theoretical analysis to formulate a set of explicit, testable hypotheses, that could be either verified or rejected in future neurophysiological experiments. Additionally, we will formulate a set of quantitive measurements to be made, to characterize a possible variability in the anisotropy or elongation of receptive fields in the primary visual cortex, with special emphasis on the relationships between a possibly predicted variability in receptive field elongation and the pinwheel structure in the primary visual cortex of higher mammals.
@@ 1.3 Contributions and novelty
  • to, in the current absence of biological measurements, about a possible variability of the degree of elongation of receptive field shapes in the primary visual cortex, provide potential indirect support for such a hypothesis, based on a combination of previously established variability in the degree of the orientation selectivity of biological receptive fields with a model-based connection between the degree of elongation of the receptive fields and their orientation selectivity, and to express possible indirect support for that hypothesis, and
  • to formulate a set of theoretically motivated and experimentally testable predictions and quantitative measurements, that could be used by experimentalists for ultimately judging whether the formulated hypothesis about a variability over the degree of elongation of the receptive fields would hold in the primary visual cortex of higher mammals, to, in turn, if that hypothesis would hold, provide potential further support for the more general hypothesis, that the receptive fields in the primary visual cortex of higher mammals may be covariant under the larger group of spatial affine transformations.
@@
  • Biological implications of that result, in that the pinwheel structure of higher mammals should, beyond an expansion over image orientations, also comprise an expansion over the degree of elongation of the receptive fields, from the center of each pinwheel to its periphery.
@@
  • The overall theoretical contributions in the paper, of explaining properties of the primary visual cortex, based on properties of a theoretically principled model for the receptive fields.
A further underlying motivation of this work is to lay out a conceptual foundation, by which theoreticians and experimentalists could join efforts, to establish to what extent the distributions of the shapes of the biological receptive fields would be compatible with an explanation from the fundamental constraint, that the family of receptive fields should be able to handle variabilities in the image data caused by geometric image transformations.
2 Methods
@@ 2.1 Affine covariant visual receptive fields
Let us represent spatial coordinates by $x=(x_{1},x_{2})^{T}$ and centered affine spatial transformations in the 2-D image domain as
@@
where $A$ represents any non-singular $2\times 2$ matrix.
Then, an affine transformed image $f^{\prime}(x^{\prime})$ of an original image $f(x)$ is defined according to
@@
With the affine transformation operator ${\cal T}_{A}$, we can write this relationship as
@@
  • applying an affine transformation $x^{\prime}=A\,x$ to an image $f(x)$ and then applying a receptive field ${\cal R}^{\prime}$ to the affine transformed image $f^{\prime}(x^{\prime})$, or
  • applying a related receptive field ${\cal R}$ to the original image $f(x)$ and then applying an affine transformation ${\cal T}_{A}$ to that output,
@@
The property of the degree of elongation of the receptive fields, to be studied in detail in this work, spans a 1-D variability within the full 4-D variability of general affine transformations, thus constituting one of the degrees of freedom in the variability of transformed receptive field shapes of ${\cal R}^{\prime}$, that will be generated by subjecting an original receptive field ${\cal R}$ to the 4-D variability of general affine transformation matrices $A$.
@@ 2.2.1 Idealized models for simple cells
  • $\varphi$ is the preferred orientation of the receptive field,
  • $\sigma_{\varphi}$ is the amount of spatial smoothing,
@@
  • $\Sigma_{\varphi}$ is a symmetric positive definite covariance matrix, with one of its eigenvectors in the direction of $\varphi$,
@@
    for $x=(x_{1},x_{2})^{T}$,
  • $\sigma_{t}$ is the amount of temporal smoothing,
  • $v=(v_{1},v_{2})^{T}$ is a local motion vector, in the direction $\varphi$ of the spatial orientation of the receptive field,
@@
  • $h(t;\;\sigma_{t})$ is a temporal Gaussian kernel with standard deviation $\sigma_{t}$.

### v10

@@ Abstract
To firmly investigate the validity of the underlying working hypothesis, we finally formulate a set of testable predictions for biological experiments, to characterize the predicted systematic variability in the elongation over the orientation maps in higher mammals, and its relations to the pinwheel structure.
@@
Receptive field Elongation Orientation selectivity Simple cell Complex cell Pinwheel
@@
Figure 2: Variability in the elongation of affine Gaussian derivative receptive fields (for the image orientation $\varphi=0$), with the scale parameter ratio $\kappa=\sigma_{2}/\sigma_{1}$ increasing from $1$ to $4\sqrt{2}$ according to a logarithmic distribution, from left to right, with the vertical scale parameter kept constant $\sigma_{2}=4$ and with the horizontal scale parameter being the smaller $\sigma_{1}\leq\sigma_{2}$. (first row) First-order directional derivatives of affine Gaussian kernels according to (5) for $m=1$. (second row) Second-order directional derivatives of affine Gaussian kernels according to (5) for $m=2$. (third row) Second-order directional derivatives of affine Gaussian kernels according to (5) for $m=3$. (fourth row) Second-order directional derivatives of affine Gaussian kernels according to (5) for $m=4$. (Horizontal axes: image coordinate $x_{1}\in[-16,16]$. Vertical axes: image coordinate $x_{2}\in[-16,16]$.)
@@ 1 Introduction
When observing objects and events in our natural environment, the image structures in the visual stimuli will be subject to substantial variabilities caused by the natural image transformations. Specifically, if observing a smooth local surface patch from different viewing directions and viewing distances, this variability can, to first order of approximation, be approximated by local affine transformations (the derivative of the projective mappings between the different views). Within the degrees of freedom of 2-D spatial affine transformations [^1] For an in-depth theoretical analysis of the degrees of freedom of 2-D spatial affine transformations in relation to the degrees of freedom of the shapes of 2-D visual receptive fields in terms of affine Gaussian derivatives, see Lindeberg (2024a). between two views of the same local surface patch, there is one degree of freedom that corresponds to non-uniform spatial scaling transformations with different amounts of scaling along two orthogonal directions. Geometrically, this degree of freedom corresponds to variabilities in the slant angle between the local surface normal and the viewing direction, as illustrated in Figure 1 with multiple views of the same scene, obtained by moving the viewing point and the viewing direction in a horizontal plane.
If we consider a vision system that views 3-D scenes from multiple viewing directions, then the receptive field responses will be strongly influenced by geometric image transformations. A particular requirement on a vision system, as imposed by the requirement of affine covariance (to be described in more detail in Sections 3.13.2), is that the vision system should be able to match the outputs from the receptive fields between different views of the same scene. Then, under the foreshortening transformations that arise from varying the slant angles under variations in the viewing conditions of 3-D objects, the spatial shapes of the receptive fields would have to be adapted to the resulting non-uniform scaling transformations, to result in different eccentricities of the receptive fields, as illustrated in Figure 2.
In Lindeberg (2021; 2023; 2024b; 2024a), we have outlined a general framework for how covariance properties with respect to geometric image transformations may constitute a fundamental constraint for the receptive fields in the primary visual cortex of higher mammals. The underlying aim is to enable the visual computations to be robust under the variabilities in the image structures generated by the natural image transformations. The fundaments of this theory are based on axiomatically determined receptive field shapes, derived from symmetry properties, that reflect structural properties of the environment, in combination with additional constraints to guarantee consistency between image representations over multiple spatial and temporal scales.
The population of receptive fields in the primary visual cortex ought to, according to this theory, obey covariance properties with respect to spatial affine transformations and Galilean transformations. Specifically, according to this theory, the receptive field shapes ought to be expanded over the degrees of freedom that correspond to the variabilities in image structures caused by these classes of geometric image transformations. Alternatively, the vision system ought to implement computational mechanisms that would to a sufficient good degree of approximation be computationally equivalent to such an expansion of receptive field shapes over the degrees of freedom of geometric image transformations. Notably, the expansion of the number of receptive fields, from about 1 M output channels from the lateral geniculate nucleus to the primary visual cortex (V1) to 190 M neurons in V1 with 37 M output channels (see DiCarlo et al. (2012) Figure 3), would indeed be consistent with an expansion of the receptive field shapes in V1 over parameters of the receptive fields.
@@
In the lack of such neurophysiological data regarding full receptive field recordings, one could, however, aim to instead obtain indirect cues regarding a possible variability in the degree of elongation of the receptive fields, by making use of the recordings of the orientation selectivity of visual neurons in previous studies: Experimental results by Nauhaus et al. (2008) show a substantial variability regarding broad vs. sharp tuning of the receptive fields in the primary visual cortex. Goris et al. (2015) report comparably uniform distributions of the degree of orientation selectivity of simple and complex cells, in terms of histograms of the resultant of the orientation selectivity curves.
@@
In a companion paper (Lindeberg 2025), we have established a strong direct link between the orientation selectivity and the elongation of the receptive fields according to the idealized generalized Gaussian derivative model for visual receptive fields (as will be summarized in Section 3.3). If the generalized Gaussian derivative shape is indeed a good model for simple and complex cells in V1, then variations in their elongation would logically correspond to a variations in their orientation selectivity.
Thereby, these results together are fully consistent with the hypothesis that the visual receptive fields in the primary visual cortex should span [^2] The terminology used throughout this paper of “spanning a variability”, is similar to the terminology in mathematics concerning a set of vectors that together span a vector space or a subspace. More generally, from the viewpoint of the wider classes of geometric image transformations in the generalized Gaussian derivative theory for visual receptive fields, one could consider these geometric image transformations to have different degrees of freedom. These degrees of freedom in the geometric image transformations will then correspond to different degrees of freedom in the shapes of the visual receptive fields, as induced by covariance properties of the receptive fields. A set of such degrees of freedom taken together will thereby span a corresponding variability in the shapes of the visual receptive fields. a variability in their anisotropy, thus consistent with the hypothesis that the receptive fields should span at least one more degree of freedom in the affine group, beyond mere rotations (as can be established by the expansion of orientation selectivity properties around the centers of the pinwheel structures in the primary visual cortex of higher mammals).
@@
To ultimately judge is this hypothesis would hold, we will finally use predictions from the presented theoretical analysis to formulate a set of explicit, testable hypotheses, that could be either verified or rejected in future neurophysiological experiments. Additionally, we will formulate a set of quantitive measurements to be made, to characterize a possible variability in the anisotropy or elongation of receptive fields in the primary visual cortex, with special emphasis on the relationships between a possibly predicted variability in receptive field elongation and the pinwheel structure in the primary visual cortex of higher mammals.
@@ 1.1 Contributions and novelty
  • In the current absence of firm biological measurements about a possible variability of the degree of elongation of receptive field shapes in the primary visual cortex, provide potential indirect support for such a hypothesis. This reasoning is based on a combination of previously established variability in the degree of the orientation selectivity of biological receptive fields with a model-based connection between the degree of elongation of the receptive fields and their orientation selectivity.
  • To formulate a set of theoretically motivated and experimentally testable predictions and quantitative measurements, that could be used by experimentalists for ultimately judging whether the formulated hypothesis about a variability over the degree of elongation of the receptive fields would hold in the primary visual cortex of higher mammals. Such a connection could then provide potential further support for the more general hypothesis, that the receptive fields in the primary visual cortex of higher mammals may be covariant under the larger group of spatial affine transformations.
@@
  • Biological implications of that result, in that the pinwheel structure of higher mammals should, beyond an expansion over image orientations, could also comprise an expansion over the degree of elongation of the receptive fields, from the center of each pinwheel to its periphery.
@@
  • The overall theoretical contributions in the paper, of explaining properties of the primary visual cortex, based on properties of a theoretically principled model for the receptive fields, using mathematical analysis and not numerical simulations of neural models as the main tool.
A further underlying motivation of this work is to lay out a conceptual foundation, by which theoreticians and experimentalists could join efforts to establish to what extent the distributions of the shapes of the biological receptive fields would be compatible with an explanation from the fundamental constraint, that the family of receptive fields should be able to handle variabilities in the image data caused by geometric image transformations.
2 Related work
Mathematical models of biological receptive fields have been formulated in terms of Gaussian derivatives (Koenderink and van Doorn 1984; 1987; 1992; Young and his co-workers 1987; 2001; 2001; Lindeberg 2013a; 2021) and Gabor filters (Marcelja 1980; Jones and Palmer 1987a; 1987b; Porat and Zeevi 1988). Gaussian derivatives have, in turn, been used as primitives in theoretical models of early visual processing by Lowe (2000), May and Georgeson (2007), Hesse and Georgeson (2005), Georgeson et al. (2007), Hansen and Neumann (2008), Wallis and Georgeson (2009), Wang and Spratling (2016), Pei et al. (2016), Ghodrati et al. (2017), Kristensen and Sandberg (2021), Abballe and Asari (2022), Ruslim et al. (2023) and Wendt and Faul (2024).
Hubel and Wiesel (1959; 1962; 1968; 2005) pioneered the study of simple and complex cells. The properties of simple cells have been further characterized by DeAngelis et al. (1995; 2004), Ringach (2002; 2004), Conway and Livingstone (2006), Johnson et al. (2008), Walker et al. (2019) and De and Horwitz (2021). Properties of complex cells have been investigated by Movshon et al. (1978), Emerson et al. (1987), Martinez and Alonso (2001), Touryan et al. (2002; 2005), Rust et al. (2005), van Kleef et al. (2010), Goris et al. (2015), Li et al. (2015) and Almasi et al. (2020), as well as modelled computationally by Adelson and Bergen (1985), Heeger (1992), Serre and Riesenhuber (2004), Einhäuser et al. (2002), Kording et al. (2004), Merolla and Boahen (2004), Berkes and Wiscott (2005), Carandini (2006), Hansard and Horaud (2011), Franciosini et al. (2019), Lindeberg (2020), Lian et al. (2021), Oleskiw et al. (2024) and Yedjour and Yedjour (2024). In this work, we will follow a specific way of modelling simple and complex cells in terms of affine Gaussian derivatives, according to the generalized affine Gaussian derivative model for visual receptive fields.
There have been previous mentions regarding receptive fields with different aspect ratios (Tinsley et al. 2003, Xu et al. 2016). It has also been reported by Wilson et al. (2016) that “neurons located near pinwheel centers … exhibit broader orientation tuning than neurons in regions of the map where neighbouring neurons exhibit similar preferences”.
According to our knowledge, there have, however, not been any previous in-depth studies in relation to a systematic expansions of receptive field shapes over the degree of elongation, nor any previously reported connections between such expansions of receptive field shapes over the degree of elongation in relation to the property of affine covariance for the family of visual receptive fields.
3 Theoretical background
@@
The material in this section will then constitute the conceptual background for the novel contributions in the paper, to be presented in Section 4, such that the paper can be read in a self-contained manner, without necessarily having to first read the references Lindeberg (2023, 2025), which this paper closely builds upon.
@@ 3.1 Affine covariant visual receptive fields
Let us represent the spatial image coordinates by $x=(x_{1},x_{2})^{T}\\ \in{\mathbb{R}}^{2}$ and centered affine spatial transformations in the 2-D image domain as
@@
where $A$ represents any non-singular $2\times 2$ matrix and $x^{\prime}=(x^{\prime}_{1},x^{\prime}_{2})^{T}\in{\mathbb{R}}^{2}$ denotes the transformed image coordinates.
Then, an affine transformed image $f^{\prime}\colon{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ of an original image $f^{\prime}\colon{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ is defined according to
@@
Let us denote the space of sufficiently smooth functions corresponding to continuous image data ${\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ by $V$. With the affine transformation operator ${\cal T}_{A}:V\rightarrow V$ from this space onto itself, we can write the affine transformation as
@@
  • applying an affine transformation $x^{\prime}=A\,x$ to an image $f$ and then applying a receptive field ${\cal R}^{\prime}:V\rightarrow V$ to the affine transformed image $f^{\prime}$, or
  • applying a related receptive field ${\cal R}:V\rightarrow V$ to the original image $f$ and then applying an affine transformation ${\cal T}_{A}$ to that output,
@@
The property of the degree of elongation of the receptive fields, to be studied in detail in this work, spans a 1-D variability within the full 4-D variability of general affine transformations. Thus, it constitutes one of the degrees of freedom in the variability of transformed receptive field shapes of ${\cal R}^{\prime}$, that will be generated by subjecting an original receptive field ${\cal R}$ to the 4-D variability of general affine transformation matrices $A$; see Lindeberg (2024a) for further details.
3.2 The hypothesis about affine covariant receptive fields
A general motivation for the wider underlying working hypothesis about affine covariance can be stated as follows: If the population of receptive fields would support affine covariance in the primary visual cortex, or sufficiently good approximations thereof, then such an ability would support the possibility of computing affine invariant image representations at higher levels in the visual hierarchy (Lindeberg 2013b). Alternatively, one could consider also sufficiently good approximations thereof, over restricted subspaces or subdomains of the most general forms of full variability under spatial affine transformations of the visual stimuli.
When we look at a slanted surface in the world, we can get a robust and stable perception of its surface texture under substantial variations of the slant angle. This robustness of visual perception under stretchings of image patterns that correspond to non-uniform scaling transformations (the perspective effects on a slanted surface patch can, to first-order of approximation, be modelled as a stretching of the image pattern along the tilt direction in image space, complemented with a uniform scaling transformation). If the visual receptive fields would span a variability under such spatial stretching transformations, then such a variability would precisely correspond to a variability in the anisotropy, or the degree of elongation, of the receptive fields.
From these theoretical motivations, one may hence raise the question whether biological vision in higher mammals would have developed computational mechanisms in the visual pathway that could be modelled in terms of affine covariance. In Lindeberg (2023, Hypothesis 1 in Section 3.2.1), it has been proposed whether the visual receptive fields in the primary visual cortex may comprise an “expansion of spatial receptive field shapes over a larger part of the affine group than mere rotations or uniform scale changes”.
The main topic of this paper is to investigate the possible validity of this hypothesis with respect to the degree of freedom of spatial affine transformations corresponding to non-uniform scaling transformations over the image domain, and corresponding to receptive field shapes of different degrees of elongation.
@@ 3.3.1 Idealized models for simple cells
  • $\varphi\in[-\pi,\pi]$ is the preferred orientation of the receptive field,
  • $\sigma_{\varphi}\in{\mathbb{R}}_{+}$ is the amount of spatial smoothing,
@@
  • $\Sigma_{\varphi}$ is a $2\times 2$ symmetric positive definite covariance matrix, with one of its eigenvectors in the direction of $\varphi$,
@@
    for $x=(x_{1},x_{2})^{T}\in{\mathbb{R}}^{2}$,
  • $\sigma_{t}\in{\mathbb{R}}_{+}$ is the amount of temporal smoothing,
  • $v=(v_{1},v_{2})^{T}\in{\mathbb{R}}^{2}$ is a local motion vector, in the direction $\varphi$ of the spatial orientation of the receptive field,
@@
  • $h(t;\;\sigma_{t})$ is a temporal Gaussian kernel with standard deviation $\sigma_{t}$ over time $t\in{\mathbb{R}}$.